OpenStudy (anonymous):

Find the complex zeros of the polynomial function f(x)=x^(3)-3^(2)+7x-5

OpenStudy (anonymous):

first, lets reduce this polynomial to a quadratic by finding one real root. you can use the ratio test.

OpenStudy (anonymous):

p : factors of 5 (the constant term) q : factors of 1 (the coefficient of x^3)

OpenStudy (primeralph):

or you can just recognize that 1+7 = 8, so a factor is (x-1)

OpenStudy (anonymous):

so, \[ p:\pm1\pm5\\ q:\pm1\\ \text{possible roots:}{p\over q}\quad \pm1,\pm5 \]

OpenStudy (anonymous):

start checking the roots using the remainder theorem: \[ x=-1\qquad f(-1)=(-1)^3-3(-1)^2+7(-1)-5=-6\ne0\implies \text{not a root}\\ f(1)=(1)^3-3(1)^2+7(1)-5=0\quad \text{YAY! root found} \]

OpenStudy (anonymous):

so, \[x=1\implies x-1\]is a factor of the polynomial. now, we use synthetic division to get the remaining quadratic factor. are you familiar with synthetic division or long division?

OpenStudy (anonymous):

I have been working on this problem for an hour, and I cant seem to get it right! I am somewhat familiar

OpenStudy (anonymous):

did you follow this first part?

OpenStudy (anonymous):

I follow you But I couldnt seem to get it myself.. Once I see it.. makes more sense!

OpenStudy (anonymous):

it is all following the steps. these are just the rules to follow through and all pieces fall into their places

OpenStudy (primeralph):

maybe you should define complexes first

OpenStudy (anonymous):

now, divide the function by this factor you can use the "draw button" below to show me your work so I can check it for ya

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

Im trying to follow! ;/

OpenStudy (anonymous):

if you followed the first part, we now have to divide. I will wait till you show the division